Teresa summerlin manufactures hand crafted pottery. Overhead is applied to products based on direct labor hours. At the beginning of the current year, Teresa estimated total overhead costs would be $30,000 and estimated direct labor hours would be 20,000. Actual overhead costs for the year totaled $36,000 for 18,000 actual direct labor hours.

By how much was overhead underapplied for the year?

A. $2,600 underapplied

B. $9,000 underapplied

C. $3,000 underapplied

D. $6,000 underapplied
Answer #1

Applied overhead = 30000*18000/20000 = $27000

Over (under) applied overhead = Applied overhead - Actual overhead

= 27000-36000

= $9000 under applied

Option B. is correct answer.